OpenStudy (anonymous):

Use the quadratic formula to find the roots of the equation. Round to tenths if necessary. 3x2 - 4x - 8 = 0 A. {-1.7, 2.8} B. {-2.8, 2} C. {-1.1, 2.4} D. no real solution

OpenStudy (rajee_sam):

do you know the quadratic formula?

OpenStudy (anonymous):

no

OpenStudy (whpalmer4):

If you have an equation in the form \(ax^2+bx+c = 0\) where \(a \ne 0\) the solutions can be found by the quadratic formula, which is \[x = \frac{-b\pm \sqrt{b^2-4ac}}{2a}\]